Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used in production (not on the basis of raw materials purchased). Its predetermined overhead rate was based on a cost formula that estimated $110,500 of manufacturing overhead for an estimated allocation base of $85,000 direct material dollars to be used in production. The company has provided the following data for the just completed year: Purchase of raw materials $ 137,000 Direct labor cost $ 81,000 Manufacturing overhead costs: Indirect labor $ 111,700 Property taxes $ 8,900 Depreciation of equipment $ 18,000 Maintenance $ 11,000 Insurance $ 11,800 Rent, building $ 32,000 Beginning Ending Raw Materials $ 23,000 $ 14,000 Work in Process $ 45,000 $ 38,000 Finished Goods $ 71,000 $ 58,000 Required:

1. Compute the predetermined overhead rate for the year.

2. Compute the amount of underapplied or overapplied overhead for the year.

3. Prepare a schedule of cost of goods manufactured for the year. Assume all raw materials are used in production as direct materials.

4. Compute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answer.

5. Assume that the $38,000 ending balance in Work in Process includes $8,600 of direct materials. Given this assumption, supply the information missing below:
Answer #1

1) Overhead rate = 110500/85000 = 130% of direct material used

2) Applied overhead = 23000+137000-14000*1.30 = 189800

Actual overhead = 193400

Under applied overhead = 189800-193400 = 3600

3) Schedule of cost of goods manufactured

Direct material used
Beginning raw material 23000
Raw material purchase 137000
Raw material available for use 160000
Less: Ending raw material -14000
Direct material used 146000
Direct labor 81000
Manufacturing overhead 189800
Total manufacturing cost 416800
Beginning work in process 45000
Total Cost of work in process 461800
Less: Ending work in process -38000
Cost of goods manufactured 423800

4) Unadjusted cost of goods sold

Beginning finished goods 71000
Cost of goods manufactured 423800
Cost of goods available for sale 494800
Less; Ending finished goods -58000
Unadjusted cost of goods sold 436800

5) Calculate following

Ending WIP 38000
Direct material 8600
Applied overhead (8600*1.30) 11180
Direct labor 18220